OpenStudy (anonymous):

if f(x)=sin^2(3-x), then f'(0)=

OpenStudy (anonymous):

Use the Chain rule and Product rule.\[\frac{ d }{ dx }\sin^{2}(3-x)*\frac{ d }{ dx }\sin(3-x)*\frac{ d }{ dx }(3-x)\]